OpenStudy (anonymous):

Evaluate the integral. 5 sec^4 x dx

OpenStudy (anonymous):

next step is... \[5 \int\limits \sec^4 x \]

OpenStudy (anonymous):

then i am lost.

OpenStudy (lgbasallote):

\[\large \int (\sec^2 x)((\sec^2 x)dx\] have you tried thaT?

OpenStudy (anonymous):

@lgbasallote so then that becomes 1/3 tan x (sec^2 x)

OpenStudy (lgbasallote):

wait is that the answer?

OpenStudy (anonymous):

i dont know the answer. i just dont know what to do after that?

OpenStudy (lgbasallote):

\[\large \int (\sec^2 x)(\sec^2 x)dx \implies \int (\tan^2 x + 1)(\sec^2 x)dx\] do you remember that?

OpenStudy (konradzuse):

ugleh!

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

what rule is that?

OpenStudy (konradzuse):

it's part of the rules of integrating sec and tan.. uploading for you.

OpenStudy (lgbasallote):

it's a trig identity \[\large \tan^2 x + 1 = \sec^2 x\]

OpenStudy (anonymous):

gotcha!

OpenStudy (lgbasallote):

so we have \[\huge \int (\tan^2 x + 1)\sec^2 x dx\] next step is to distribute sec^2 x \[\huge \int (\tan^2 x \sec^2 x + \sec^2 x)dx\]

OpenStudy (konradzuse):

I have circled what's important for this section.

OpenStudy (anonymous):

so how do you integrate that?

OpenStudy (lgbasallote):

you distribute the integral symbol \[\int \tan^2 x \sec^2 xdx + \int \sec^2 x dx\] the second term is easy to integrate..integrate that first

OpenStudy (anonymous):

that would become....tan x?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

so the whole answer would be 5 tan x +5/x tan^3 x +c

OpenStudy (anonymous):

\[ \tan^3 x \over 3 \]

OpenStudy (lgbasallote):

then the first term is u-sub

OpenStudy (lgbasallote):

wait what did you do?

OpenStudy (anonymous):

well the whole problem was 5 sec^4 x dx

OpenStudy (anonymous):

we were just doing the sec^4 x

OpenStudy (anonymous):

part

OpenStudy (lgbasallote):

lol i know i just cant understand your notation sorry \[\Large 5[\int \tan^2 x \sec^2 x dx + \int sec^2 x dx]\] \[\Large 5[\frac{\tan^3 x}{3} + \tan x]\] \[\Large \frac{5\tan^3 x}{3} + 5\tan x\] is this what you meant?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

well then you're right! :D